# Singleton algorithms for the Constraint Satisfaction Problem

### v1

@@ 4 Limitations of linear programming
Horn-SAT $\mathrm{AIP}$ $\mathrm{ArcCons}$
2-SAT $\mathrm{BLP}$, $\mathrm{Singl}\mathrm{AIP}$ $\mathrm{Singl}\mathrm{ArcCons}$, $\mathrm{ArcCons}+\mathrm{AIP}$
@@
$\left(\begin{smallmatrix}0&1&2&3&4\\ 1&0&3&4&2\end{smallmatrix}\right)$ $\mathrm{BLP}+\mathrm{AIP}$ $\mathrm{Singl}\mathrm{ArcCons}$, $\mathrm{Singl}\mathrm{AIP}$
@@
$\mathbb{D}_{4}$ $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ ???
@@ 4.2 Disconnected templates
$\left(\begin{smallmatrix}\mathbf{0}&\mathbf{1}&\mathbf{2}&\mathbf{3}&\mathbf{4}\\ \mathbf{1}&\mathbf{0}&\mathbf{3}&\mathbf{4}&\mathbf{2}\end{smallmatrix}\right)=(\{0,1,2,3,4\};\{(0,1),(1,0),(2,3),(3,4),(4,2)\})$, which is a structure containing a single binary relation given by the graph of a bijection. To show that both $\mathrm{Singl}\mathrm{ArcCons}$ and $\mathrm{Singl}\mathrm{AIP}$ solve it, let us define $f\in\operatorname{Pol}\left(\begin{smallmatrix}0&1&2&3&4\\ 1&0&3&4&2\end{smallmatrix}\right)$ that is simultaneously $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block totally symmetric and palette block alternating. Specifically, let $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ return the value from the first constant block if it exists, and $\mathbf{x}_{1}^{1}$ otherwise.
@@
Thus, we have linear equations modulo 2 on $\{0,1\}$ and independent linear equations modulo 3 on $\{0^{\prime},1^{\prime},2^{\prime}\}$. Any instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$ decomposes into two independent systems, one over $\mathbb{Z}_{2}$ and one over $\mathbb{Z}_{3}$. Solving the instance amounts to solving each system separately and returning “Yes” if both have solutions. Nevertheless, $\mathrm{AIP}$ fails here: it returns “Yes” on any instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$. The key idea comes from the following claim (the definition of a $d$-solution was given in Section 3.3).
@@
In contrast, $\mathrm{Singl}\mathrm{AIP}$ solves the problem. Fixing a concrete element in the $\mathrm{Singl}$-algorithm restricts the instance to either the $\{0,1\}$-part or the $\{0^{\prime},1^{\prime},2^{\prime}\}$-part, and each part can be solved by $\mathrm{AIP}$. This is also witnessed by a $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block alternating function $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ defined as follows:
@@ 4.3 Implications of linear equations
To witness that $\mathrm{Singl}\mathrm{AIP}$ correctly solves $\operatorname{CSP}(\mathbb{Z}_{2}\leftarrow 2)$ we define a $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette block alternating function $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ from $\operatorname{Pol}(\mathbb{Z}_{2}\leftarrow 2)$ as follows:
@@ 4.4 Dihedral group
Suppose $n,\ell\geq 4$. Then $\operatorname{Clo}(x\circ y)$ does not contain an idempotent $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric operation.
@@
5 Dihedral group $\mathbf{D}_{4}$.
Claim 5.1.
Proof.
Similarly, we derive
$f^{3}((0,0,x_{1}^{3}),\dots,(0,0,x_{n}^{3}))=$
Claim 5.2.
$f^{1}(x_{1},\dots,x_{n})$ $=g(x_{1}^{1},\dots,x_{n}^{1}),$
$f^{2}(x_{1},\dots,x_{n})$ $=g(x_{1}^{2},\dots,x_{n}^{2}).$
Proof.
Claim 5.3.
Proof.
Claim 5.4.
$f^{1}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{1}+\dots+a_{n}x_{n}^{1},$
$f^{2}(x_{1},\dots,x_{n})$ $=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2},$
Claim 5.5.
$\mathcal{D}_{4}\subseteq\operatorname{Clo}(x\circ y)$.
Proof.
Claim 5.6.
$\operatorname{Pol}(\mathbb{D}_{4})\subseteq\mathcal{D}_{4}$.
Proof.
for all $b_{1},\dots,b_{n},d_{1},\dots,d_{n}\in\mathbb{Z}_{2}$. These two conditions imply that $g(z_{1},\dots,z_{n})=a_{1}z_{1}+\dots+a_{n}z_{n}$ and $g_{3}$ returns 0 whenever the first $n$ or the last $n$ arguments equal 0. Therefore, the polynomial representation of $g_{3}$ cannot have a monomial $x_{i}^{1}$ or $x_{i}^{2}$. By Claims 5.3 and 5.4, the only possible monomials are of the form $x_{i}^{1}\cdot x_{j}^{2}$. Let $g_{3}(x_{1}^{1},\dots,x_{n}^{1},x_{1}^{2},\dots,x_{n}^{2})=\sum\limits_{i,j\in[n]}c_{i,j}x_{i}^{1}x_{j}^{2}$. It remains to check that the relation $R$ enforces the property $c_{i,j}+c_{j,i}=a_{i}\cdot a_{j}$. Substituting the tuple $((1,0,0),(0,1,0))$ into the $i$-th position, $((0,1,0),(1,0,0))$ into the $j$-th position, and $((0,0,0),(0,0,0))$ into the remaining positions of $f$, we obtain the tuple $((a_{i},a_{j},c_{i,j}),(a_{j},a_{i},c_{j,i}))$, which must belong to $R$. Hence, the definition of $R$ gives $a_{i}\cdot a_{j}+c_{i,j}+c_{j,i}=0$. ∎
Lemma 4.1.
Proof.
Lemma 4.3.
Suppose $n,\ell\geq 4$. Then $\operatorname{Clo}(x\circ y)$ does not contain an idempotent $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric operation.
Proof.
Assume the contrary. Let $f\in\operatorname{Clo}(x\circ y)$ be such an operation. By Lemma 5.6, $f\in\mathcal{D}_{4}$, so we can choose $a:[n]\to\mathbb{Z}_{2}$ and $c:[n]\times[n]\to\mathbb{Z}_{2}$ defining $f$. Take $i,j\in[\ell+1]$ with $i\neq j$. Substitute $(1,0,0)$ for the $i$-th coordinate and into the blocks 3 and 4 (the second overlined part), substitute $(0,1,0)$ for the $j$-th coordinate and into the blocks 5 and 6 (the third overlined part), and $(0,0,0)$ for the remaining coordinates. Since $f$ is $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette block symmetric, the result must be the same for all $i$ and $j$. The first coordinate of the result is $a_{i}+a_{2\ell+2}+a_{2\ell+3}+\dots+a_{4\ell+2}$, and the third one is
@@ 6 Characterization of Singleton Algorithms
In this section we prove the characterization of the singleton algorithms in terms of palette block symmetric functions. We start with a formal definition of a minion and show how abstract minions can be used to characterize the power of algorithms. Then, in Section 6.2 we apply the Hales-Jewett theorem to show that, under certain conditions, a homomorphic image of a minion contains more symmetric objects than the original. In the last two subsections we show how to apply this result to obtain a characterization of the algorithms $\mathrm{Singl}\mathrm{ArcCons}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, respectively.
@@ 6.3 Characterization of the singleton arc-consistency algorithm
Let $\mathcal{M}_{\mathrm{SArc}}$ be the minion whose N{N}-ary part consists of matrices $M\colon N\times[L]\to\{0,1\}$, where $L\in\mathbb{N}$, such that:
@@
  • $\mathrm{CSingl}\mbox{{ArcCons}}$ returns “Yes” on $\mathbb{X}$;
@@
$\Rightarrow$ Assume that $\mathrm{CSingl}\mbox{{ArcCons}}$ returns “Yes” on $\mathbb{X}$. First, notice that finding a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$ is equivalent to finding matrices from $\mathcal{M}_{\mathrm{SArc}}$ for each constraint and each variable such that they are compatible. Since the algorithm $\mathrm{CSingl}\mbox{{ArcCons}}$ returns “Yes”, for every constraint $C=R(x_{1},\dots,x_{s})$ we obtain a nonempty relation $R_{f}$ (the final version of $R$ in the pseudocode). Let $P_{1},\dots,P_{L}$ be the list of all pairs $(C,\mathbf{a})$, where $C$ is a constraint and $\mathbf{a}$ is a tuple from the corresponding $R_{f}$. Then for every $i\in[L]$, the algorithm $\mathrm{ArcCons}$ returns a nonempty subset $D_{y}^{i}$ for each variable $y$. The matrix $M_{x}\colon A\times[L]\to\{0,1\}$ corresponding to the variable $x$ is defined by $M_{x}(a,i):=\begin{cases}1,&\text{if $a\in D_{x}^{i}$}\\ 0,&\text{otherwise}\end{cases}$. Similarly, the matrix $M_{C}\colon R^{\mathbb{A}}\times[L]\to\{0,1\}$ corresponding to a constraint $C=R(x_{1},\dots,x_{s})$ is defined by
@@
It follows from arc-consistency that this assignment gives us a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$.
$\Leftarrow$ Assume that $h\colon\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$ is a homomorphism. First, notice that it is sufficient to prove the claim for connected instances $\mathbb{X}$, since the algorithm $\mathrm{CSingl}\mbox{{ArcCons}}$ essentially works with disconnected parts independently. Thus, we assume that $\mathbb{X}$ is connected. Then for any two variables $x_{1},x_{2}\in X$, the matrices $h(x_{1})$ and $h(x_{2})$ have the same width $L\in\mathbb{N}$. For each constraint $C=R(x_{1},\dots,x_{s})$, let $R^{\prime}$ be the set of all tuples $\mathbf{a}\in R$ such that the $\mathbf{a}$-th row of the matrix corresponding to $C$ has a non-zero entry. Since the matrices in $\mathcal{M}_{\mathrm{SArc}}$ are skeleton, for each tuple $\mathbf{a}\in R^{\prime}$ there exists $\ell_{C,\mathbf{a}}\in[L]$ such that the $\ell_{C,\mathbf{a}}$-th column of the matrix corresponding to $C$ has 1 only in the $\mathbf{a}$-th position. Let $D_{y}^{C,\mathbf{a}}$ be the set of all $b\in D_{y}$ such that $h(y)(b,\ell_{C,\mathbf{a}})=1$. Since the sets $(D_{y}^{C,\mathbf{a}}\colon y\in X)$ form an arc-consistent reduction (see Observation 3.1), the arc-consistency algorithm must return “Yes” after we restrict the constraint $C$ to $\mathbf{a}$. Hence, $\mathrm{CSingl}\mathrm{ArcCons}$ returns “Yes”. ∎
@@
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature. Then $\mathrm{CSingl}\mathrm{ArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if there exists a minion homomorphism $\mathcal{M}_{\mathrm{SArc}}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
@@
Suppose $n,\ell\in\mathbb{N}$, $\mathcal{N}$ is a minion whose $n$-ary part is finite, $h\colon\mathcal{M}_{SArc}\to\mathcal{N}$ is a minion homomorphism. Then $h(\mathcal{M}_{SArc})$ contains an $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette block totally symmetric object.
@@
Let us check the required conditions (1)-(3) of Theorem 6.4. Condition (1) follows from the definition of $\phi$. For condition (2), let $\sigma\in\Omega$. For each $i$ appearing in $\sigma$ (viewed as a tuple), let $u(i)$ be the position of a block that contains only $i$ (such a block exists since $\sigma$ is palette). Then $\phi(u(i))^{\sigma}$ has $1$ at the $i$-th position and 0 elsewhere. Hence, $(\phi(u(\sigma(1)))\circ\dots\circ\phi(u(\sigma(N))))^{\sigma}=I^{\sigma}$. Condition (3) holds because adding $I$ makes the matrix skeleton. By Theorem 6.4 we obtain $M\in\mathcal{M}_{\mathrm{SArc}}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$. Thus, $h(M)$ is an $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette block totally symmetric object. ∎
@@
Suppose $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette block totally symmetric functions for all $\ell_{1},\dots,\ell_{n}\in\mathbb{N}$. Then $\mathrm{SinglArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
@@
Notice that by identifying variables within block of an $(L_{1},\dots,L_{n})$-palette block totally symmetric (PBTS) function we obtain a PBTS function with smaller block sizes. Therefore, for all $L_{1},\dots,L_{n}\in\mathbb{N}$ we can choose an $(L_{1},\dots,L_{n})$-PBTS function $f_{L_{1},\dots,L_{n}}\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$ such that identifying variables within blocks of $f_{L_{1},\dots,L_{n}}$ yields $f_{\ell_{1},\dots,\ell_{n}}$, where $\ell_{i}\leq L_{i}$ for every $i$. To simplify notation, we write $f(S_{1},\dots,S_{n})$ meaning $f_{k_{1},\dots,k_{n}}(s_{1}^{1},\dots,s_{1}^{k_{1}},\dots,s_{n}^{1},\dots,s_{n}^{k_{n}})$, where $S_{i}=\{s_{i}^{1},\dots,s_{i}^{k_{i}}\}$ for every $i\in[n]$. This notation is justified because we only apply $f_{k_{1},\dots,k_{n}}$ to palette tuples, so the order of elements in the sets $S_{i}$ does not matter.
@@
    $\mathrm{CSingl}\mathrm{ArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
@@
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette block totally symmetric function for every $n,m_{1},\dots,m_{n}\in\mathbb{N}$;
@@
    for every $N\in\mathbb{N}$ there exist $n,m_{1},\dots,m_{n}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette block totally symmetric function.

### v2

@@ 4 Limitations of linear programming
Horn-SAT $\mathrm{AIP}$ $\mathrm{AC}$
2-SAT $\mathrm{BLP}$, $\mathrm{Singl}\mathrm{AIP}$ $\mathrm{Singl}\mathrm{AC}$, $\mathrm{BLP}+\mathrm{AIP}$
@@
$\left(\begin{smallmatrix}0&1&2&3&4\\ 1&0&3&4&2\end{smallmatrix}\right)$ $\mathrm{BLP}+\mathrm{AIP}$ $\mathrm{Singl}\mathrm{AC}$, $\mathrm{Singl}\mathrm{AIP}$
@@
$\mathbb{D}_{4}^{idemp}$ $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$ ??? [^1] We already know that $\operatorname{CSP}(\mathbb{D}_{4}^{idemp})$ can be solved by the second level of $\mathrm{AIP}$; however higher-level algorithms are outside the scope of this paper.
@@ 4.2 Disconnected templates
$\left(\begin{smallmatrix}\mathbf{0}&\mathbf{1}&\mathbf{2}&\mathbf{3}&\mathbf{4}\\ \mathbf{1}&\mathbf{0}&\mathbf{3}&\mathbf{4}&\mathbf{2}\end{smallmatrix}\right)=(\{0,1,2,3,4\};\{(0,1),(1,0),(2,3),(3,4),(4,2)\})$, which is a structure containing a single binary relation given by the graph of a bijection. To show that both $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}\mathrm{AIP}$ solve it, let us define $f\in\operatorname{Pol}\left(\begin{smallmatrix}0&1&2&3&4\\ 1&0&3&4&2\end{smallmatrix}\right)$ that is simultaneously $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette totally symmetric and palette alternating. Specifically, let $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ return the value from the first constant block if it exists, and $\mathbf{x}_{1}^{1}$ otherwise.
@@
Thus, we have linear equations modulo 2 on $\{0,1\}$ and independent linear equations modulo 3 on $\{0^{\prime},1^{\prime},2^{\prime}\}$. Any instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$ decomposes into two independent systems, one over $\mathbb{Z}_{2}$ and one over $\mathbb{Z}_{3}$. Solving the instance amounts to solving each system separately and returning “Yes” if one of them has a solution. Nevertheless, $\mathrm{AIP}$ fails here: it returns “Yes” on any instance of $\operatorname{CSP}(\mathbb{Z}_{2}\cup\mathbb{Z}_{3})$. The key idea comes from the following claim (the definition of a $d$-solution was given in Section 3.3).
@@
In contrast, $\mathrm{Singl}\mathrm{AIP}$ solves the problem. Fixing a concrete element in the $\mathrm{Singl}$-algorithm restricts the instance to either the $\{0,1\}$-part or the $\{0^{\prime},1^{\prime},2^{\prime}\}$-part, and each part can be solved by $\mathrm{AIP}$. This is also witnessed by a $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette alternating function $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ defined as follows:
@@ 4.3 Implications of linear equations
To witness that $\mathrm{Singl}\mathrm{AIP}$ correctly solves $\operatorname{CSP}(\mathbb{Z}_{2}\leftarrow 2)$ we define a $(\underbrace{\overline{2\ell+1},\dots,\overline{2\ell+1}}_{n})$-palette alternating function $f(\mathbf{x}_{1},\dots,\mathbf{x}_{n})$ from $\operatorname{Pol}(\mathbb{Z}_{2}\leftarrow 2)$ as follows:
@@ 4.4 Dihedral group
Suppose $n,\ell\geq 4$. Then $\operatorname{Clo}(x\circ y)$ does not contain an idempotent $(\underbrace{\overline{\ell+1,\ell},\overline{\ell+1,\ell},\dots,\overline{\ell+1,\ell}}_{2n})$-palette symmetric operation.
@@ 5 Characterization of Singleton Algorithms
In this section we prove the characterization of the singleton algorithms in terms of palette symmetric functions. We start with a formal definition of a minion and show how abstract minions can be used to characterize the power of algorithms. Then, in Section 5.2 we apply the Hales-Jewett theorem to show that, under certain conditions, a homomorphic image of a minion contains more symmetric objects than the original. In the last two subsections we show how to apply this result to obtain a characterization of the algorithms $\mathrm{Singl}\mathrm{AC}$ and $\mathrm{Singl}(\mathrm{BLP}+\mathrm{AIP})$, respectively.
@@ 5.3 Characterization of the singleton arc-consistency algorithm
We characterize minions for the singleton versions via matrices, following the ideas of Ciardo and Živný in [23]. Let $\mathcal{M}_{\mathrm{SArc}}$ be the minion whose N{N}-ary part consists of matrices $M\colon N\times[L]\to\{0,1\}$, where $L\in\mathbb{N}$, such that:
@@
  • $\mathrm{CSingl}\mbox{{AC}}$ returns “Yes” on $\mathbb{X}$;
@@
$\Rightarrow$ Assume that $\mathrm{CSingl}\mbox{{AC}}$ returns “Yes” on $\mathbb{X}$. First, notice that finding a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$ is equivalent to finding matrices from $\mathcal{M}_{\mathrm{SArc}}$ for each constraint and each variable such that they are compatible. Since the algorithm $\mathrm{CSingl}\mbox{{AC}}$ returns “Yes”, for every constraint $C=R(x_{1},\dots,x_{s})$ we obtain a nonempty relation $R_{f}$ (the final version of $R$ in the pseudocode). Let $P_{1},\dots,P_{L}$ be the list of all pairs $(C,\mathbf{a})$, where $C$ is a constraint and $\mathbf{a}$ is a tuple from the corresponding $R_{f}$. Then for every $i\in[L]$, the algorithm $\mathrm{AC}$ returns a nonempty subset $D_{y}^{i}$ for each variable $y$. The matrix $M_{x}\colon A\times[L]\to\{0,1\}$ corresponding to the variable $x$ is defined by $M_{x}(a,i):=\begin{cases}1,&\text{if $a\in D_{x}^{i}$}\\ 0,&\text{otherwise}\end{cases}$. Similarly, the matrix $M_{C}\colon R^{\mathbb{A}}\times[L]\to\{0,1\}$ corresponding to a constraint $C=R(x_{1},\dots,x_{s})$ is defined by
@@
Let us show that the matrices for constraints and for variables are consistent, that is, for any constraint $C=R(x_{1},\dots,x_{s})$ and $i\in[s]$, we have $M_{C}^{\pi_{i}}=M_{x}$, where $\pi_{i}\colon R^{\mathbb{A}}\to A$ is a projection onto the $i$-th coordinate. Since the $i$-th columns of all matrices come from the output of the arc-consistency algorithm, the $i$-th columns are consistent. Taking a minor of a matrix object is equivalent to taking the corresponding minor of each column and then gluing the results, so the matrices themselves are consistent. Hence the assignment we defined yields a homomorphism $\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$.
$\Leftarrow$ Assume that $h\colon\mathbb{X}\to\mathbb{F}_{\mathcal{M}_{\mathrm{SArc}}}(\mathbb{A})$ is a homomorphism. First, notice that it is sufficient to prove the claim for connected instances $\mathbb{X}$, since the algorithm $\mathrm{CSingl}\mbox{{AC}}$ essentially works with disconnected parts independently. Thus, we assume that $\mathbb{X}$ is connected. Then for any two variables $x_{1},x_{2}\in X$, the matrices $h(x_{1})$ and $h(x_{2})$ have the same width $L\in\mathbb{N}$. For each constraint $C=R(x_{1},\dots,x_{s})$, let $R^{\prime}$ be the set of all tuples $\mathbf{a}\in R$ such that the $\mathbf{a}$-th row of the matrix corresponding to $C$ has a non-zero entry. Since the matrices in $\mathcal{M}_{\mathrm{SArc}}$ are skeleton, for each tuple $\mathbf{a}\in R^{\prime}$ there exists $\ell_{C,\mathbf{a}}\in[L]$ such that the $\ell_{C,\mathbf{a}}$-th column of the matrix corresponding to $C$ has 1 only in the $\mathbf{a}$-th position. Let $D_{y}^{C,\mathbf{a}}$ be the set of all $b\in D_{y}$ such that $h(y)(b,\ell_{C,\mathbf{a}})=1$. Since the sets $(D_{y}^{C,\mathbf{a}}\colon y\in X)$ form an arc-consistent reduction (see Observation 3.1), the arc-consistency algorithm must return “Yes” after we restrict the constraint $C$ to $\mathbf{a}$. Hence, $\mathrm{CSingl}\mathrm{AC}$ returns “Yes”. ∎
@@
Suppose $\mathbb{A}$ and $\mathbb{B}$ are finite relational structures in the same signature. Then $\mathrm{CSingl}\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$ if and only if there exists a minion homomorphism $\mathcal{M}_{\mathrm{SArc}}\to\operatorname{Pol}(\mathbb{A},\mathbb{B})$.
@@
Suppose $n,\ell\in\mathbb{N}$, $\mathcal{N}$ is a minion whose $n$-ary part is finite, $h\colon\mathcal{M}_{SArc}\to\mathcal{N}$ is a minion homomorphism. Then $h(\mathcal{M}_{SArc})$ contains an $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette totally symmetric object.
@@
Let us check the required conditions (1)-(3) of Theorem 5.4. Condition (1) follows from the definition of $\phi$. For condition (2), let $\sigma\in\Omega$. For each $i$ appearing in $\sigma$ (viewed as a tuple), let $u(i)$ be the position of a block that contains only $i$ (such a block exists since $\sigma$ is palette). Then $\phi(u(i))^{\sigma}$ has $1$ at the $i$-th position and 0 elsewhere. Hence, $(\phi(u(\sigma(1)))\circ\dots\circ\phi(u(\sigma(N))))^{\sigma}=I^{\sigma}$. Condition (3) holds because adding $I$ makes the matrix skeleton. By Theorem 5.4 we obtain $M\in\mathcal{M}_{\mathrm{SArc}}^{(N)}$ such that $h(M)^{\sigma_{1}}=h(M)^{\sigma_{2}}$ for any $(\sigma_{1},\sigma_{2})\in\omega$. Thus, $h(M)$ is an $(\underbrace{\overline{\ell},\dots,\overline{\ell}}_{n})$-palette totally symmetric object. ∎
@@
Suppose $\mathbb{A}$ and $\mathbb{B}$ are relational structures in the same signature, $\mathbb{A}$ is finite, $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has $(\overline{\ell_{1}},\dots,\overline{\ell_{n}})$-palette totally symmetric functions for all $\ell_{1},\dots,\ell_{n}\in\mathbb{N}$. Then $\mathrm{SinglArcCons}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$.
@@
Notice that by identifying variables within blocks of an $(L_{1},\dots,L_{n})$-palette totally symmetric (PTS) function we obtain a PTS function with smaller block sizes. Therefore, for all $L_{1},\dots,L_{n}\in\mathbb{N}$ we can choose an $(L_{1},\dots,L_{n})$-PTS function $f_{L_{1},\dots,L_{n}}\in\operatorname{Pol}(\mathbb{A},\mathbb{B})$ such that identifying variables within blocks of $f_{L_{1},\dots,L_{n}}$ yields $f_{\ell_{1},\dots,\ell_{n}}$, where $\ell_{i}\leq L_{i}$ for every $i$. To simplify notation, we write $f(S_{1},\dots,S_{n})$ meaning $f_{k_{1},\dots,k_{n}}(s_{1}^{1},\dots,s_{1}^{k_{1}},\dots,s_{n}^{1},\dots,s_{n}^{k_{n}})$, where $S_{i}=\{s_{i}^{1},\dots,s_{i}^{k_{i}}\}$ for every $i\in[n]$. This notation is justified because we only apply $f_{k_{1},\dots,k_{n}}$ to palette tuples, so the order of elements in the sets $S_{i}$ does not matter.
@@
    $\mathrm{CSingl}\mathrm{AC}$ solves $\operatorname{PCSP}(\mathbb{A},\mathbb{B})$;
@@
    $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette totally symmetric function for every $n,m_{1},\dots,m_{n}\in\mathbb{N}$;
@@
    for every $N\in\mathbb{N}$ there exist $n,m_{1},\dots,m_{n}\geq N$ such that $\operatorname{Pol}(\mathbb{A},\mathbb{B})$ has an $(\overline{m_{1}},\dots,\overline{m_{n}})$-palette totally symmetric function.